### v2

@@ 3.1 The Population-Level Testing Problem
A disadvantage of RUM bound is that it requires data on $\mathcal{D}/X$ to be consistent with RUM. There are two possible issues with this. First, it could be that the population choice distributions on on $\mathcal{D}/X$ are inconsistent with RUM; in that case, $p_{D}^{RUM}(X)$ is not well-defined. Second, it could be that rationalizable population distributions generate non-rationalizable finite sample frequencies; in that case, one could define a feasible version of $p_{d}^{RUM}(X)$, and we will do so in Section 4.3.
It remains to clarify how we can test stochastic rationalizability of data. In its essence, this question has been resolved in McFadden and Richter (1991). Because we adapt their approach to an environment in which rationalizability is not characterized by standard revealed preference axioms (although they mention the possibility of this generalization) and to keep the treatment self-contained, we will briefly explain the answer. The basic insight is that probabilistic choice data can be rationalized if, and only if, it can be expressed as convex combination of data that would be produced by deterministic utility maximizers. [^11] As is often the case with seminal results, with ample hindsight the intuition informs a much easier than the original proof (Stoye, 2019).
@@
To make this precise, construct a matrix $\bm{A}$ s.t. each row of $\bm{A}$ corresponds to a given alternative within a particular choice set (i.e., each alternative appears once for every choice set containing it), and each rationalizable deterministic choice pattern corresponds to exactly one column of $\bm{A}$. For example, consider a data set consisting of choices from $\left\{a_{1},d\right\}$, $\left\{a_{1},a_{2},d\right\}$ and $\left\{a_{1},a_{2},a_{3},d\right\}$. The first row of $\bm{A}$ indicates the choice of $a_{1}$ from $\{a_{1},d\}$, the second row the choice of $d$ from $\{a_{1},d\}$ and so on. One column of $\bm{A}$ (namely, the first one in 3.1 below) then represents the choices of someone who picked $a_{1}$ from all three choice sets, which can be rationalized as the choices of a decision maker who prefers $a_{1}$ over all other alternatives in $X$. The remaining columns of $\bm{A}$ represent all other rationalizable choice patterns. Some book-keeping reveals that in this example, one has: [^12] The size of $\bm{A}$ increases extremely rapidly with the complexity of data; see Kitamura and Stoye (2018) for numerical examples.
$\begin{array}[]{c}a_{1}|\left\{a_{1},d\right\}\\ d|\left\{a_{1},d\right\}\\ a_{1}|\left\{a_{1},a_{2},d\right\}\\ a_{2}|\left\{a_{1},a_{2},d\right\}\\ d|\left\{a_{1},a_{2},d\right\}\\ a_{1}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{2}|\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{3}|\left\{a_{1},a_{2},a_{3},d\right\}\\ d|\left\{a_{1},a_{2},a_{3},d\right\}\end{array}\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&0&0&0&0&0&0\\ 0&0&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}=\bm{A}.$ (3.1)
The core insight of McFadden and Richter (1991) is summarized in the following theorem.
@@
Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{A}$ derived from some stochastic choice data set. That data set is rationalizable by RUM if and only if there exists a vector $\nu\in\Delta^{H-1}$ (the $(H-1)$-dimensional unit simplex, where $H$ is the number of columns of $\bm{A}$) such that
$\bm{A}\nu=\pi$
In order to adapt this approach to our idealized data set, we need to deal with the fact that we only observe whether or not the default was chosen. To do so, we premultiply the matrix $\bm{A}$ with a matrix $\bm{B}$ that merges events we do not separately observe – i.e. the choice of different non-default options. In the above example, we would have
$\bm{B}=\left\{\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&1&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right\}$
A corollary to Theorem 3.1 then characterizes $p_{d}^{RUM}(X)$:
@@
Let $\pi$ be the observed choice probabilities associated with each row of the matrix $\bm{BA}$ derived from a data set $\{\mathcal{D},p\}$ which reports only the probability of default choice. That data set is rationalizable by RUM if and only if their exists a probability vector $\nu$ such that
$\bm{BA}\nu=\pi.$
@@
For intuition, observe that in (3.1), the vector $\bm{a}$ is the last row of $\bm{A}$. Equivalently, it is the indicator of the choice type being the one whose choice is always the default (this is the only type who chooses the default from $X$). The bound simply maximizes the probability corresponding to this type, subject to overall data being stochastically rationalizable.
In practice, $p_{d}^{RUM}(X)$ as stated is well-defined only if $\pi$ is stochastically rationalizable. Since the set of rationalizable probability vectors is “small” (we will elaborate on this in Section 4.3), empirical choice frequencies will typically –and do in our data– fail this. In that case, a feasible version of the bound can be computed by substituting a constrained estimator of $\pi$. This will be illustrated later.
We finally note that the above machinery opens the door to using this testing approach on other, for instance non-RUM, models. Nothing in the above development forces $\bm{A}$ to contain columns that correspond exactly to conventionally rationalizable behaviors. By adding (removing) columns of $\bm{A}$, one can test less (more) restrictive models. Indeed, in our empirical application we will use this approach to test:
@@
    A model which admits all the rationalizable deterministic choice patterns of RUM but also allows for choosing $d$ from choice set $X$ (regardless of behavior on smaller sets).
@@
    The same model as (ii), except that a choice type may also switch to choosing $d$ for all choice sets of cardinality $3$. (Types that choose $d$ from all sets of cardinality $3$ must also choose $d$ from choice set $X$.)
Models (ii) and (iii) capture ‘choice overloaded’ behavior in that the only violations of standard rationality they allow are switches to the default when the choice set expands. Moving from (i) to (iii) enlarges $\bm{A}$, and therefore increases computational cost, but does not add conceptual difficulties. Since the resulting models are nested (listed in increasing order of generality), one can then ask what is the least permissive model that is not rejected in the data. This will allow us to unpack what sort of choice overloaded behavior, if any, could have generated our data; for reasons that will become clear, this line of analysis will be central to our results. As a preview, model (iii) will be the only one not rejected in our data. This analysis is in Section 4.3, where we also argue that model (iii) is itself a restrictive model.
We note that a further generalization which captures choice overload would be to allow switches to the default not as a function of set cardinality (as in (ii) and (iii)), but along any distinct path of choice set expansion. Dean (2008) considers a similar model. We do not consider this generalization due to computational complexity (the corresponding $\bm{A}$-matrix is very large) and because we know it would not be rejected in our data given it nests (iii).
@@ 3.2 Econometric Tests
We next explain some testing strategies that connect the above ideas to recent advances in econometrics. To this purpose, we consider samples that were generated by randomly drawing individuals and then exposing each individual to an (i.i.d. randomly generated) selection of choice problems, i.e. subsets of $\mathcal{D}$ [^13] This mirrors our empirical design, which was partly chosen because, unlike stratified sampling or mean-reverting coins, it is easy to bootstrap. In particular, data may contain choices from the same individuals in different choice sets, as is the case in our application. Note also that in this section, the term “data” is used in the statistical sense, i.e. referring to a finite sample from an underlying population.
We begin with two tests of the Min bound. We will estimate choice probabilities $p_{d}(\cdot)$ by the analogous sample frequencies, which can be informally defined as
$\hat{p}_{d}(A)\equiv\frac{\sum\bm{1}\{\text{choice problem is }A,\text{ choicemade is }d\}}{\sum\bm{1}\{\text{choice problem is }A\}},$ (3.3)
where $\bm{1}\{\cdot\}$ is the indicator function and the sums are taken over all data points. For all but the finite-sample test that we present first, any estimator of $p_{d}(\cdot)$ whose asymptotic distribution is normal or approximated by the simple nonparametric bootstrap would suffice. Sample averages have both properties as long as they are not close to degenerate; by direct inspection, this condition is easily met in our data.
@@ 3.2.1 A Finite-Sample Test of the Min Bound
The second expression clarifies that this is a joint test of potentially many hypotheses. Consider first testing any one of these, i.e. testing whether $p_{d}(X)\leq p_{d}(A)$ for a specific $A\subset X$. To this purpose, define $\hat{p}_{d,A}(X)$ in exact analogy to (3.3) but dropping observations from subjects who also saw choice problem $A$. In words, we estimate $p_{d}(X)$ by its simple empirical analog among subjects who did not see $A$. As a result, $\hat{p}_{d}(A)$ and $\hat{p}_{d,A}(X)$ estimate binomial proportions in two mutually exclusive samples. We will therefore apply Fisher’s (1992) exact test for binomial proportions to $H_{0}:p_{d}(X)\leq p_{d}(A)$ for any given $A$.
Of course, we need to account for the fact that we will conduct many such tests once ($78$ in our application) and cannot assume independence. Our first approach is Bonferroni adjustment, that is, all p-values are multiplied by $78$. An advantage of this approach is that it ensures finite-sample (as opposed to asymptotic) size control. However, its power is limited through three channels: The estimators $\hat{p}_{d,A}(\cdot)$ discard data; Fisher’s exact test is in general conservative due to integer issues; and Bonferroni adjustment is conservative. In practice, with the sample sizes that we generated for our empirical application, we expect only the last channel to have an appreciable effect.
@@ 3.2.2 An Asymptotic Test of the Min Bound
The finite sample test adjusts for the fact that, in principle, many tests are conducted simultaneously. A common concern with such adjustments is that, if the results of some of these tests appear obvious, one might needlessly lose power. Indeed, consider the visualization of our data in Figure 3: The default probabilities in the right-hand cluster are obviously much higher than in the grand choice set (corresponding to the dashed line). Can we restrict attention to only those inequality conditions that might reasonably bind?
This question has received considerable attention in the econometric literature on moment inequalities and is by now well understood. We implement a method that can be seen as special case of Andrews and Soares (2010) and also of Chernozhukov et al. (2013), both of whom establish its validity under rather general conditions.
The method can in principle use many test statistics; for concreteness, set
@@
The test will reject if $t$ is too large. The catch is that the distribution of $t$, and therefore the appropriate critical value, depends on the nuisance parameter $(p_{d}(X)-p_{d}(A))_{A\in\mathcal{D}\setminus X}$. This parameter cannot be pre-estimated with sufficient accuracy [^14] Technically, it enters the asymptotic distribution scaled by $\sqrt{n}$. Our exposition is slightly informal because the issue is well understood in the literature; see also Canay and Shaikh (2017, section 4) for a survey. and so we must conservatively approximate it. This is done in three steps:
@@
    where $\hat{p}_{d}^{*}(\cdot)$ denotes the bootstrap analog of $\hat{p}_{d}(\cdot)$. This bootstrap will be clustered by individual, i.e. we (i.i.d. uniformly with replacement) resample individuals and use all responses from a given resampled individual; this ensures that correlation patterns in $(\hat{\pi}-\pi)$ due to eliciting several responses per individual are captured.
@@
    Use a pre-test with size converging to 00, e.g. $\alpha_{n}=1/\log(n)$, and discard from consideration any sets $A$ s.t. the null hypothesis $H_{0}:p_{d}(X)\geq p_{d}(A)$ is rejected at significance level $\alpha_{n}$. Let $\mathcal{D}^{*}$ denote the set of choice problems that are retained in this pre-test.
@@
This procedure reflects two important ideas from the moment inequalities literature. First, the bootstrap population of data must be on the null hypothesis, which necessitates a recentering. In our case, the least favorable and therefore relevant instance of the null hypothesis is that all relevant probabilities are equal. Since the test statistic is location invariant, for concise notation and implementation we recenter them to 00. This is reflected in the definition of $\tilde{p}_{d}(\cdot)$. Second, the test may be extremely conservative if we accordingly recenter all $78$ estimators. Therefore, we pre-screen choice items whose default probability is likely to much exceed $p_{d}(X)$. This is a special case of Generalized Moment Selection (Andrews and Soares, 2010). In particular, the size of the pre-test must go to 00 in order for us to claim that we will asymptotically select all binding constraints. [^15] Our depiction of the method is simplified by picking a specific test statistic and also filling in specific values for several tuning parameters. For example, the sum $\sum_{A\in\mathcal{D}^{*}\setminus X}\max\{\hat{p}_{d}(X)-\hat{p}_{d}(A),0\}$ of constraint violations might yield a test statistic that is more powerful against small but uniform violations. Also, rather than letting the size of a pre-test vanish, one could use Bonferroni correction to “spend” some “coverage budget” on the pre-test (Andrews and Barwick, 2012; Romano et al., 2014). Again, we omit these generalizations for conciseness because they are known in the literature. The test that we implement is as stated, and the rejection in our empirical application is so resounding that none of these choices would plausibly have impacted it.

### v3

@@ 3.1 The Population-Level Testing Problem
A feature of the RUM bound is that it requires choice from $\mathcal{D}/X$ to be consistent with RUM. There are two possible issues with this. First, it could fail at the population level; in that case, $p_{D}^{RUM}(X)$ is not well-defined. Second, rationalizable population distributions will still, at least occasionally, generate non-rationalizable finite sample frequencies; in that case, one could define a feasible version of $p_{d}^{RUM}(X)$, and we will do so in Section 4.2.
Notice that this feature of the RUM bound, along with our focus only on default choice probabilities, means that many well known behavioral phenomena will not show up as choice overload as per our definition – either because they would lead to violations of RUM in the set $\mathcal{D}/X$, or because they would not be observable due to the coarsening of our data. For example, consider a case in which the DM was first asked to choose between two hard to compare alternatives – $x$ and $y$, say – plus a default, and then asked to choose from the same set with an alternative $z$ that is dominated by $x$ but not $y$. The asymmetric dominance effect might push people to switch their choices from $y$ to $x$ in the larger choice set, but if this does not affect the choice of default, it will not show up as choice overload.
It remains to clarify how we can test stochastic rationalizability of data. For the case of standard stochastic choice data, this question has been resolved by McFadden and Richter (1991, see () for a short proof). Here we adapt their approach to our data set.
@@
The basic insight is that choice probabilities can be rationalized if, and only if, they can be expressed as convex combination of data that would be produced by deterministic utility maximizers. To make this precise, construct a matrix $\bm{A}$ s.t. each row of $\bm{A}$ corresponds to a given alternative within a particular choice set (i.e., each alternative appears once for every choice set containing it). Each column corresponds to a deterministic choice pattern rationalizable by a different (strict) preference profile over $X$. For example, let $\mathcal{D}=\{\{a_{1},d\},\{a_{1},a_{2},d\},\{a_{1},a_{2},a_{3},d\}\}$ and let the first row of $\bm{A}$ indicate choice of $a_{1}$ from $\{a_{1},d\}$, the second row choice of $d$ from $\{a_{1},d\}$, and so on. One column of $\bm{A}$ (namely, the first one in (3.1)) then represents the choices of someone who picked $a_{1}$ from all three choice sets, rationalizable by preferences ranking $a_{1}$ first. Constructing the remaining columns from all other rationalizable choice patterns, one has:
$\begin{array}[]{c}a_{1}\mid\left\{a_{1},d\right\}\\ d\mid\left\{a_{1},d\right\}\\ a_{1}\mid\left\{a_{1},a_{2},d\right\}\\ a_{2}\mid\left\{a_{1},a_{2},d\right\}\\ d\mid\left\{a_{1},a_{2},d\right\}\\ a_{1}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{2}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ a_{3}\mid\left\{a_{1},a_{2},a_{3},d\right\}\\ d\mid\left\{a_{1},a_{2},a_{3},d\right\}\end{array}\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&0&0&0&0&0&0\\ 0&0&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}=\bm{A}.$ (3.1)
McFadden and Richter’s (1991) core insight is the following.
@@
Let the vector $\rho$ collect observed choice probabilities in order corresponding to the rows of $\bm{A}$. These probabilities are rationalizable by RUM if, and only if, there exists a vector $\nu\in\Delta^{H-1}$ (the $(H-1)$-dimensional unit simplex, where $H$ is the number of columns of $\bm{A}$) such that
$\bm{A}\nu=\rho.$
We slightly adapt this approach because we assume that we only observe whether $d$ was chosen from any choice set. Thus, premultiply $\bm{A}$ by a matrix $\bm{B}$ that merges the choice of different non-default options in a given menu. Columns of $\bm{B}\bm{A}$ then represent different rationalizable deterministic default vs non-default choice patterns. Each choice set in $\mathcal{D}$ is represented by two rows in $\bm{B}\bm{A}$, the first of which indicates non-default choice from the choice set, and the second representing default choice. For the example above, $\bm{B}$ and $\bm{BA}$ are as follows:
$\bm{B}=\left\{\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&1&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right\},\quad\bm{BA}=\left\{\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&0&0&0&1&1&1&1\\ 1&1&1&1&0&0&1&1\\ 0&0&0&0&1&1&0&0\\ 1&1&1&1&0&1&1&1\\ 0&0&0&0&1&0&0&0\end{array}\right\}$
A corollary to Theorem 3.1 then characterizes $p_{d}^{RUM}(X)$. To state it formally, let the vector $\pi$ collect probabilities of active and passive choice in order corresponding to rows of $\bm{BA}$; in our example,
$\pi=\begin{pmatrix}1-p_{d}(\{a_{1},d\})\\ p_{d}(\{a_{1},d\})\\ 1-p_{d}(\{a_{1},a_{2},d\})\\ p_{d}(\{a_{1},a_{2},d\})\\ 1-p_{d}(\{a_{1},a_{2},a_{3},d\})\\ p_{d}(\{a_{1},a_{2},a_{3},d\})\end{pmatrix}.$
Then we can write:
@@
A probability vector $\pi$ as just defined is rationalizable by RUM if, and only if, there exists $\nu\in\Delta^{H-1}$ such that
$\bm{BA}\nu=\pi.$ (3.2)
@@
For intuition, observe that in (3.1), the vector $\bm{a}$ is the last row of $\bm{A}$. Equivalently, it is the indicator of the unique choice type whose choice is always the default. The bound simply maximizes the probability of this type, subject to overall data being stochastically rationalizable. [^6] Expression (3.3) presupposes that $\pi$ is stochastically rationalizable. Since the set of rationalizable probability vectors is “small” (we will elaborate on this in Section 4.2), empirical choice frequencies may fail this. In that case, a feasible version of the bound can be computed by substituting a constrained estimator of $\pi$. This will be illustrated later.
Note that, when we only observe default vs non-default choice, multiple deterministic rational choice types may be indistinguishable – e.g., the first four columns of $\bm{A}$ represent types that never choose the default from choice problems in $\mathcal{D}$ and hence are identical in $\bm{BA}$. In practice, one may simplify problem (3.2) by eliminating such repetitions.
We finally note that we can apply this testing approach to other (e.g. non-RUM) models. Nothing forces $\bm{A}$ to contain columns that correspond exactly to conventionally rationalizable behaviors. By adding (removing) columns of $\bm{A}$, one can test less (more) restrictive models.
We will use this observation as the basis of method for identifying choice overload. We will empirically test:
@@
    A relaxation of RUM that allows for choosing $d$ from choice set $X$, regardless of behavior on smaller sets.
@@
    The same model as (ii), except that a choice may switch to $d$ for all choice sets of cardinality $3$. (Types that choose $d$ from all sets of cardinality $3$ must also choose $d$ from $X$.)
The RUM bound is violated if (i) is rejected while model (ii) is not: Model (ii) captures choice overloaded behavior by only allowing violations of standard rationality through switching to $d$ when the choice set expands.
Model (iii) expands on the standard notion of choice overload by allowing it to “kick in” at smaller choice sets. Moving from (i) to (iii) enlarges $\bm{A}$ but does not add conceptual difficulties. Since the resulting models are nested, one can then ask what is the least permissive model that is not rejected in the data. This will allow us to unpack what sort of choice overloaded behavior, if any, could have generated our data. In Section 4.2, we show that model (iii) is the only one not rejected in our data and argue that is still a restrictive model.
@@ 3.2 Econometric Tests
We next explain testing strategies that connect the above ideas to recent advances in econometrics. [^7] This section can be skipped without loss of continuity. To this purpose, we consider samples that were generated by randomly drawing individuals and then giving each individual a (i.i.d. randomly generated) selection of choice problems. [^8] This mirrors our empirical design, which was partly chosen because, unlike stratified sampling or mean-reverting coins, it is easy to bootstrap. In particular, data may contain choices from the same individuals in different choice sets, as is the case in our application.
We estimate choice probabilities $p_{d}(\cdot)$ by the analogous sample frequencies $\hat{p}_{d}(A)$. For all but the finite sample test that we present first, any estimator of $p_{d}(\cdot)$ whose asymptotic distribution is normal or approximated by the simple nonparametric bootstrap would suffice. Sample frequencies have both properties as long as they are not close to degenerate, where “close to” is relative to sample size. We chose a relatively large sample size precisely to ensure this, and it easily holds in our data.
@@ 3.2.1 A Finite-Sample Test of the Min Bound
The second expression clarifies that this is a joint test of potentially many hypotheses. Consider first testing any one of these, i.e. testing whether $p_{d}(X)\leq p_{d}(A)$ for a specific $A\subset X$. To this purpose, define the “leave-$A$-out” sample frequency $\hat{p}_{d,A}(X)$ by dropping observations from subjects who also saw choice problem $A$. As a result, $\hat{p}_{d}(A)$ and $\hat{p}_{d,A}(X)$ estimate binomial proportions in two mutually exclusive samples. We therefore apply Fisher’s (1992) exact test for binomial proportions to $H_{0}:p_{d}(X)\leq p_{d}(A)$ for any given $A$.
Of course, we need to account for the fact that we conduct many such tests once ($78$ in our application) and cannot assume independence. Our first approach is Bonferroni adjustment, that is, all p-values are multiplied by $78$. An advantage of this approach is that it ensures finite-sample (as opposed to asymptotic) size control. However, its power is limited through three channels: The estimators $\hat{p}_{d,A}(\cdot)$ discard data; Fisher’s exact test is in general conservative due to integer issues and a strong sense of conditional validity; Bonferroni adjustment is conservative. In practice, with the sample sizes that we generated for our empirical application, we expect only the last channel to have an appreciable effect.
@@ 3.2.2 An Asymptotic Test of the Min Bound
The finite sample test adjusts for the fact that, in principle, many tests are conducted simultaneously. A common concern with such adjustments is that, if the results of some of these tests appear obvious, one might needlessly lose power. Indeed, in our experimental data, the default probabilities in a number of choice sets are obviously much higher than in the grand choice set. [^9] Figure 7 in Supplemental Appendix A.2 provides a visualization of this. Can we restrict attention to only those inequality conditions that might reasonably bind?
This question has received considerable attention in the econometric literature on moment inequalities. We implement a method that can be seen as special case of Andrews and Soares (2010) and also of Chernozhukov et al. (2013), both of whom establish its validity under rather general conditions.
The method can use many test statistics; for concreteness, set
@@
The test will reject if $t$ is too large. The catch is that the distribution of $t$, and therefore the appropriate critical value, depends on the nuisance parameter $(p_{d}(X)-p_{d}(A))_{A\in\mathcal{D}\setminus X}$. This parameter cannot be pre-estimated with sufficient accuracy [^10] Technically, it enters the asymptotic distribution scaled by $\sqrt{n}$. See Canay and Shaikh (2017, section 4) for a survey. and so we must conservatively approximate it. This is done in three steps:
@@
    where $\hat{p}_{d}^{*}(\cdot)$ denotes the bootstrap analog of $\hat{p}_{d}(\cdot)$. This bootstrap will be clustered by individual, i.e. we (i.i.d. uniformly with replacement) resample individuals and use all responses from a given resampled individual; this ensures that correlation patterns in $\hat{p}_{d}(\cdot)$ due to eliciting several responses per individual are captured.
@@
    Use a pre-test with size converging to 00, e.g. $\alpha_{n}=\alpha/\log(n)$, where $\alpha$ is the test’s nominal size. Discard from consideration any sets $A$ s.t. the null hypothesis $H_{0}:p_{d}(X)\geq p_{d}(A)$ is rejected at significance level $\alpha_{n}$. Let $\mathcal{D}^{*}$ denote the set of choice problems that are retained in this pre-test.
@@
This procedure reflects two important ideas from the moment inequalities literature. First, the bootstrap population of data must be on the null hypothesis, which necessitates a recentering. In our case, the least favorable and therefore relevant instance of the null hypothesis is that all relevant probabilities are equal. Since the test statistic is location invariant, for concise notation and implementation we recenter them to 00. This is reflected in the definition of $\tilde{p}_{d}(\cdot)$. Second, the test may be extremely conservative if we accordingly recenter all $78$ estimators. Therefore, we pre-screen choice items whose default probability is likely to much exceed $p_{d}(X)$.In particular, because the size of the pre-test goes to 00, we will asymptotically select all binding constraints. [^11] Our depiction of the method is simplified by picking a specific test statistic and also filling in specific values for several tuning parameters. Also, rather than letting the size of a pre-test vanish, one could use Bonferroni correction to “spend” some “coverage budget” on the pre-test (Andrews and Barwick, 2012; Romano et al., 2014). This would make no difference in our empirical application because its p-values are far from conventional thresholds.